OpenStudy (anonymous):

What is the simplified form of the reciprocal of √-48 ?

OpenStudy (anonymous):

1/4sqrt(3)i

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{1}{\sqrt{-48}}=\frac{1}{\sqrt{16\times 3\times -1}}\] \[=\frac{1}{4\times \sqrt{3}i}\] \[=\frac{1}{4\times \sqrt{3}i}\times \frac{\sqrt{3}i}{\sqrt{3}i}\] \[=-\frac{\sqrt{3}}{12}i\]